Find the pull on the plunger of an electromagnet when the total flux uniformly distributed is 0.8 mWb. The diameter of the plunger is 5 cm.


1
Expert's answer
2022-03-28T14:28:40-0400

Answer


That pulling force is given by

"F=\frac{\phi^2}{2A\mu_0}\\=\frac{0.8*0.8*10^{-6}}{2*3.14*0.05*0.05*4*3.14*10^{-7}}\\=129.69N"